Documentation for fancybox.sty: Box tips and tricks for LaTEX

Timothy Van Zandt tvz@Princeton.EDU Version 1.0 February 10, 1993

fancybox.sty, together with its documentation, gives extensive answers to and solutions for many questions about how to frame or rotate this or that in LaTEX. It also contains commands for shadow, double and oval frames.

Contents

1 Fancy frames 1 2 A short course on boxes 2 3 Defining LR-box environments 3 4 Math 5 5 Floats 7 6 Center, flushleft and flushright 10 7 Lists 12 8 Superimposing boxes 14 9 Framing a whole page 15 10 Switching to landscape mode midstream 17 11 Verbatim 18 12 Changes 27

1 Fancy frames

fancybox.sty has five variants of LaTEX’s \fbox command:

\shadowbox, \doublebox, \ovalbox (with \thinlines) and
\Ovalbox (with \thicklines).

Here are examples:^{1}

\shadowbox{\large\bf New Glarus Birdwatch}

New Glarus Birdwatch

\doublebox{\large\bf New Glarus Birdwatch}

New Glarus Birdwatch

\ovalbox{\large\bf New Glarus Birdwatch}

[figure]

New Glarus Birdwatch

\Ovalbox{\large\bf New Glarus Birdwatch}

[figure]

New Glarus Birdwatch

The distance between the box and the frame is \fboxsep, as with LaTEX’s \fbox command. The commands use other parameters as well:

\shadowbox The width of the frame is \fboxrule (the same as with \fbox). The width of the shadow is \shadowsize (default: 4pt).

\doublebox The width of the inner frame is .75\fboxrule, and the width of the outer frame is 1.5\fboxrule. The distance between the two frames is 1.5\fboxrule plus .5pt.

1In this documentation, the default value of \fboxsep has been changed from 3pt to 6pt.

\ovalbox The width of the frame is set by the \thinlines declaration. The diameter of the corner arcs is set with the \cornersize command.

         \cornersize{num}

sets the diameter of the corners arcs to num times the lessor of the width and height of the box.

         \cornersize*{dim}

sets the diameter of the corner arcs to dim. This is all approximate, because LaTEX has a limited range of arc sizes to choose from. The default is

         \cornersize{.5}

\Ovalbox This is like \ovalbox, except that the width of the lines is set by the \thicklines declaration.

There are no analogs to LaTEX’s \framebox command, which has various optional arguments not supported by \fbox. You can get the exact same functionality by putting the argument of the above framing commands in a

\makebox.

There is also a variant \fancyoval of LaTEX’s \oval picture object. The difference is that \oval always makes the diameter of the corner arcs as large as possible, and \fancyoval uses the \cornersize command to set the diameter.

2 A short course on boxes

The \shadowbox, \doublebox, \ovalbox and \Ovalbox commands described

in the previous section are examples of LR-box commands, meaning that their argument is processed in LR mode. LaTEX LR-box commands include

\mbox, \makebox, \fbox, \framebox, \sbox and \savebox. All the PSTricks

commands whose argument is text are LR-box commands, including, e.g, the framing, rotating, scaling and positioning commands, and some of the node commands. Any rotation command is an LR-box command.

The purpose of the rest of this documentation is to provide answers to, and solutions for, frequently asked questions about using LR-box commands with LaTEX. I will use \fbox for the leading example of a box framing command,^{2} and \rotateleft for the leading example of a box rotation command. (fancybox.sty does not contain a \rotateleft command, as this must be implemented via \special’s, but there are numerous box-rotation style files around.) However, most of what is said here applies to any LR-box command.

In each LR-box command, the text is processed in restricted horizontal mode, which is referred to as “LR-mode” in Lamport’s LaTEX: User’s Guide and Reference Manual. In restricted horizontal mode, the input, consisting of regular characters and boxes, is made into one (long or short) line. There is no line-breaking, nor can there be vertical mode material such as an entire displayed equation. However, the fact that you can include another box means that this isn’t really a restriction.

For one thing, alignment environments such as LaTEX’s tabular are just boxes, and thus present no problem. Picture environments and the LR-box commands themselves are also just boxes. Entire paragraphs or other vertical mode material such as displayed equations can be nested in a \parbox or

minipage.

3 Defining LR-box environments

To frame a minipage, you have to write

     \fbox{%
       \begin{minipage}{3in}
         blah
       \end{minipage}}

You might want to define an environment fminipage that frames its contents, but you can’t use

     \newenvironment{fminipage}%
       {\fbox{\begin{minipage}}%
       {\end{minipage}}}

because the braces are not balanced in the definition. fancybox.sty contains an Sbox environment that makes it easy to define your own LR-box environments. It is a variant of LaTEX’s \sbox command that saves the contents of the environment in a storage bin that can be retrieved with the command \TheSbox.^{3} For example, here is a framed minipage:

2In the examples using \fbox, be aware that the default value of \fboxsep has been changed in this documentation from 3pt to 6pt.

     \begin{Sbox}
       \begin{minipage}{3in}
         blah
       \end{minipage}
     \end{Sbox}
     \fbox{\TheSbox}

and here is an fminipage environment that works:

     \newenvironment{fminipage}%
       {\begin{Sbox}\begin{minipage}}%
       {\end{minipage}\end{Sbox}\fbox{\TheSbox}}

Let’s see that it really works:

     \begin{fminipage}{2in}
       Since the former doesn't use braces to delimit
       the contents of the box, $\ldots$
     \end{fminipage}

Since the former doesn’t use braces to delimit the contents of the box, …

      \begin{Sbox}
        blah
      \end{Sbox}
      \TheSbox

and

3The difference between

\newsavebox{\mybox}

\sbox{\mybox}{blah}

\usebox{\mybox}

is that Sbox saves the contents globally, and \TheSbox erases the contents globally.

4 Math

In-line math, or pieces of a displayed equation (as opposed to a whole equation), are horizontal mode material, but most LR-box commands switch out of math mode when they occur in math mode. Thus, you have to explicitly switch back in to math mode when desired.^{4} For example:

     $x + y = \fbox{$\Omega$}$

x + y = Ω

You also have to explicitly write

     \scriptstyle, \scriptscriptstyle or \displaystyle

if you want one of these special math styles. For example, here I will frame an equation, but not the equation number:

     \begin{equation}
       \fbox{$\displaystyle
         \int_{\Omega_0} \zeta(\omega) d\omega
         \geq \bar{r}$}
     \end{equation}

∫_{Ω_{0}} ζ(ω)dω ≥ r̄  (1)

Entire displayed equations or eqnarray environments work differently because they are vertical mode material. Thus, they have to go inside a \parbox

or minipage. E.g.,
     \newlength{\mylength}
     \[
       \setlength{\fboxsep}{15pt}
       \setlength{\mylength}{\linewidth}
       \addtolength{\mylength}{-2\fboxsep}

4This is not true for the PSTricks LR-box commands.

       \addtolength{\mylength}{-2\fboxrule}
       \fbox{%
         \parbox{\mylength}{
           \setlength{\abovedisplayskip}{0pt}
           \setlength{\belowdisplayskip}{0pt}
           \begin{equation}
             x + y = z
           \end{equation}}}
     \]

x + y = z (2)

The outer \[ \] are just used to display the boxed equation, rather than actually switch into math mode. Note how I set the width of the \parbox so that the displayed box would exactly have width \linewidth.^{5}

I also set the display skips to 0pt and increased the size of \fboxsep so that I would have the same distance all around between the equation and the frame.

This is again a mouthful, and so I might instead define:^{6}

     \newenvironment{FramedEqn}%
       {\setlength{\fboxsep}{15pt}
         \setlength{\mylength}{\linewidth}%
         \addtolength{\mylength}{-2\fboxsep}%
         \addtolength{\mylength}{-2\fboxrule}%
         \Sbox
         \minipage{\mylength}%
            \setlength{\abovedisplayskip}{0pt}%
            \setlength{\belowdisplayskip}{0pt}%
            \equation}%
       {\endequation\endminipage\endSbox
         \[\fbox{\TheSbox}\]}

5That is what \mylength is for. It is better to define a single scratch length that you reuse rather than creating a new one each time.

6The reason for using \minipage instead of \begin{minipage}, and so on, is that with AmS-LaTEX, \begin and \end cannot appear in the definition of a new equation environment.

fancybox.sty doesn’t bother defining any such environments, because there are too many possible designs. But let’s see if the one above works:

     \begin{FramedEqn}
       \Rightarrow P\sim\xi(P_\gamma)- \frac{1}{3}
     \end{FramedEqn}

⇒ P ∼ ξ(P_{γ}) − \frac{1}{3} (3)

fancybox.sty contains a Beqnarray environment, which is like the eqnarray environment but it is not vertical mode material. Instead, it produces a box just large enough to hold all the equations. For example:

     \fbox{%
       \begin{Beqnarray*}
         x & = & y\\
         y & > & x \\
         \int_4^5 f(x)dx & = & \sum_{i\in F} x_i
       \end{Beqnarray*}}

\begin{aligned}x & = & y \\ y & > & x \\ ∫_{4}^{5} f(x)dx & = & ∑_{i∈F} x_{i} \end{aligned}

The unstarred version produces standard equation numbers on the right (even with the leqno style option and AmS-LaTEX). It might not work with special equation numbering macros.

5 Floats

A common mistake is to put a whole table, figure or other float environment inside an LR-box command. Instead, you should put everything that is inside the environment (including the \caption, if you want that boxed too) inside a minipage of the desired width, and then put the minipage inside the LR-box command.

For example:

     \begin{table}[h]
       \begin{center}
         \fbox{%
           \begin{minipage}{.8\textwidth}
             \begin{center}
              \begin{tabular}{rl}
                foo & bar
              \end{tabular}
             \end{center}
             \caption{A table of foo and bar.}
           \end{minipage}}
       \end{center}
     \end{table}

foo bar

Table 1: A table of foo and bar.

Note how I had to use center twice: once to center the framed box, and again to center the stuff inside the box.

That is a mouthful, and so I might define a FramedTable environment like the following, which sets the size of the minipage so that the framed box is exactly the width of the page (no need for the first center environment this time):

     \newenvironment{FramedTable}%
       {\begin{table}[h]
         \begin{Sbox}%
         \setlength{\mylength}{\textwidth}%
         \addtolength{\mylength}{-2\fboxsep}%
         \addtolength{\mylength}{-2\fboxrule}%

         \begin{minipage}{\mylength}}%
       {\end{minipage}\end{Sbox}\fbox{\TheSbox}\end{table}}%

Now let’s see if it works:

     \begin{FramedTable}
       \begin{center}
         \begin{tabular}{rl}
           foo & bar
         \end{tabular}
       \end{center}
       \caption{A table of foo and bar.}
     \end{FramedTable}

foo bar

Table 2: A table of foo and bar.

The most common reason to want to rotate an entire float, caption and all, is to put it on a page by itself in landscape mode, centered both horizontally and vertically. Compared to the table framing we did above, we just have to replace \fbox by our box rotation command (e.g., \rotateleft or whatever), set the width of the minipage to \textheight (if you want to use the full size of the page), and use the float position specifier [p]. fancybox.sty contains an environment,

     \begin{landfloat}{float}{rotation command}
       ...
     \end{landfloat}

that automates this. It has two arguments: the name of the floating environment, and your rotation command. For example, if \rotateleft{foo} rotates foo by 90 degrees, and you want a landscape mode table, then try

     \begin{landfloat}{table}{\rotateleft}
       ...
     \end{landfloat}

If the whole document is in landscape mode, then landfloat gives you a portrait-mode float—good for a table that is too tall to fit in landscape mode.

If you don’t add a caption to a float, it doesn’t matter much what floating environment you use (e.g., table, figure or whatever). Thus, you can put anything in a landscape float. For example, suppose I have a very wide equation. Then I can write:

     \begin{landfloat}{table}{\rotateleft}
       \begin{equation}
         ...
       \end{equation}
     \end{landfloat}

fancybox.sty defines a generic caption, \GenericCaption, that doesn’t affect the numbering of floats, doesn’t make an entry in a list of floats, and doesn’t add anything to the argument you give. I could have used this if I wanted to add a caption to the previous example.

6 Center, flushleft and flushright

There are two ways to box a center, flushleft or flushright environment. If you have long lines and you want TEX to do the line breaking, then put the environment inside a minipage.^{7} If you have short lines and you want the frame to adjust itself to the size of the longest line, then use fancybox.sty’s

     Bcenter, Bflushleft or Bflushright

environment instead.

These are basically just tabular environments with a single column, and so each line should end with \\, or \\[dim] to insert extra space, and the text on each line must be balanced. Also, each line is sitting in its own group, and so, e.g., if you want to change the font for the whole environment, you should do this before the environment rather than after. Like the tabular environment, the box that is produced has the baseline at the center, unless you include an optional argument [t] to align the box with the top line or [b] to align the box with the bottom line.

7List and other such environments work best inside a minipage environment rather than a \parbox.

For example:

     \setlength{\fboxsep}{10pt}%
     \fbox{%
       \begin{Bcenter}
         Love of life\\
         and other short stories\\
         by Policarpa Salabarrieta
     \end{Bcenter}}

Love of life and other short stories by Policarpa Salabarrieta

Compare this with:

     \setlength{\fboxsep}{10pt}%
     \fbox{%
       \begin{minipage}{6cm}
       \begin{center}
         Love of life and other short stories
         by Policarpa Salabarrieta
       \end{center}
       \end{minipage}}

Love of life and other short stories by Policarpa Salabarrieta

In either case, if we want the resulting framed box centered, we have to include it inside another center environment.

Here is another example:

     My list: \fbox{%
       \begin{Bflushleft}[b]

         Galanga root\\
         Coconut\\
         Tempeh
       \end{Bflushleft}}

My list: Galanga root Coconut Tempeh

7 Lists

There are again two ways to box a list environment such as

     itemize, enumerate or description.

You can put it in a minipage of pre-determined size and let TEX do the line breaking, or you can use

     Bitemize, Benumerate or Bdescription

instead, and let the box adjust to the size of the longest line.

For example.

     \fbox{%
       \begin{Bitemize}
         \item Groceries
         \item Clean hamster cages
         \item Pick up Peter
       \end{Bitemize}}

• Groceries
• Clean hamster cages
• Pick up Peter

Most of the usual list parameters are irrelevant except for \itemsep and \labelsep. These environments are also based on the tabular environment, and so each item should be balanced text. You can’t write \vspace{4pt} either, but you can insert an extra amount of space before an item by writing \item(4pt) (or \item(4pt)[label] if you have a label). Also, you can break lines within an item using \\ or \\[dim]. For example:

     \fbox{%
       \begin{Bdescription}
         \item[David] Groceries
         \item[Eli]  Hamster cages\\
                    Surreal numbers
         \item(3pt)[Doris] Pick up Peter
       \end{Bdescription}}

David Groceries
Eli  Hamster cages

Surreal numbers

Doris Pick up Peter

These environments also have an optional argument, [t] to align the box with the top line, and [b] to align the box with the bottom line.

     To do:
     \fbox{\setlength{\itemsep}{0pt}%
       \begin{Benumerate}[t]
         \item Groceries
         \item Hamster cages
         \item Pick up Peter
       \end{Benumerate}}

To do: 1. Groceries 2. Hamster cages 3. Pick up Peter

There is also a generic \Blist command that is analogous to LaTEX’s \list. It has the same two obligatory arguments, plus a third optional [t] or [b] argument for changing the alignment.

8 Superimposing boxes

The command

\boxput*(x,y){LR stuff1 }{LR stuff2 }

puts LR stuff1 either behind (the default) or in front of (with the *) LR stuff2 .^{8} The resulting box has the dimensions of LR stuff2 .

The coordinates (x,y) determine where the center of LR stuff1 is positioned. For example, (0,0) puts it at the center of LR stuff2 , (0,1) puts it at the center-top, and (-1,-1) puts it in the bottom-left corner.

More generally, the origin of the coordinate system is at the center of LR stuff2 , one unit in the vertical direction is half the vertical size of LR stuff2 , and one unit in the horizontal direction is half the width of LR stuff2 . Thus, x and y should always be numbers (without units such as pt or cm), with one exception: If y is b or B, LR stuff1 is positioned vertically at the baseline of LR stuff2 . (x,y) is optional—the default is (0,0).

Except for the funny coordinate system, \boxput is like the \put command in a picture. In particular, you can use \makebox in LR stuff1 to fine-tune the positioning, and LR-stuff1 can contain a picture environment.

You might use \boxput to put a “water mark” in the background of a box, or to put a label next to a box, when you don’t want this label to take up any space. Here is a lazy example:

     \boxput{\makebox(0,0){\Huge Censored!}}{\parbox{3in}{%
       The origin of the coordinate system is at the center of
       {\em LR stuff2}, and one unit in the x-direction
       is half the width of {\em LR stuff2}.}}

Censored!

The origin of the coordinate system is at the center of LR stuff2, and one unit in the x-direction is half the width of LR stuff2.

This would be a lot more interesting using PSTricks, with “Censored!” in the foreground, rotated 30 degrees, and red:

8You will only notice the difference between \boxput and \boxput* if you are using color implemented by \special’s

     \boxput*{\rput{30}{\Huge\red Censored}}{\parbox{3in}{%
       blah blah}}

Here is another example using PSTricks:

     \newcommand{\titledframe}[2]{%
       \boxput*(0,1){\psframebox*{#1}}%
         {\psframebox[framesep=12pt]{#2}}}

The following example illustrated roughly how it works, but “My title” does not blot out the frame behind it because this documentation does not use PSTricks.

     \titledframe{My title}{%
       \parbox{2in}{The title is superimposed
         on the top part of the frame.}}

The title is superimposed on the top part of the frame.

My title

9 Framing a whole page

The commands

     \fancyput*(x,y){LR stuff }
     \thisfancyput*(x,y){LR stuff }

are pretty much like \put commands in a LaTEX picture environment whose origin is 1 inch down and to the right from the top left corner of the page.^{9} The only differences are that (i) that any LR-mode material is permitted (including LaTEX picture environment, of course), (ii) the coordinate is optional (0pt,0pt) is substituted by default), and (iii) if the coordinate is included, you must specify the units.

\thisfancyput affects only the current page, and is a global declaration

(analogous to \thispagestyle).

9Don’t blame me for TEX’s peculiar 1 inch margins.

[figure]

If you include the optional *, then the command adds to, rather than replaces, other things that have been inserted with \fancyput or \thisfancyput.

These commands are particularly useful for framing a page, because you can get a frame that is, e.g., 1 inch from each side of the physical page without having to worry about what margins you are using for the document. Here is an example:

     \thisfancyput(3.25in,-4.5in){%
       \setlength{\unitlength}{1in}\fancyoval(7,9.5)}

You could also use \fancyput to add some kind of “watermark” or background image (e.g., a light gray “DRAFT”).

There are other commands that directly frame or in some other way box the page of text:

     \fancypage{cmds1 }{cmds2 }
     \thisfancypage{cmds1 }{cmds2 }

Each finished page, before adding the headers and footers, (and thus having width and height \textwidth and \textheight, is boxed with

cmds1 {pagebox}

Thus, cmds1 should be, or should end with, a command whose argument can be a box, such as \fbox or \rotateleft.

Then the headers and footers are added, using the new width of the page, and this is boxed with

cmds2 {pagebox}

The same rules apply to cmds2 as to cmds1 .

Here is an example:

     \thisfancypage{%
       \setlength{\fboxsep}{8pt}%
       \setlength{\shadowsize}{8pt}%
       \shadowbox}{}

Warning: The commands described in this section change LaTEX’s output routine, and may not work with document styles that do the same. Also, bad arguments can cause serious errors with uninformative error messages.

10 Switching to landscape mode midstream

The most common reason to switch to landscape mode midstream is to rotate a float, such as a table or figure. This was discussed in Section 5.

If you want to rotate one or more pages, without rotating the headers and footers, use the

     \begin{LandScape}{cmd}
       ...
     \end{LandScape}

environment. cmd should be the command for rotating the page 90 degrees to the left or right. (E.g., \rotateleft, or \rotate[l].)

If you want to rotate the headers, footers and margins as well, use the

\begin{Landscape}{paperwidth}{paperheight}{cmd}

       ...
     \end{Landscape}

environment (the small s makes the difference) to rotate the pages left (counterclockwise), and use the Landscape* environment (same arguments) to rotate the pages right (clockwise). The three arguments are the width of the paper, the height of the paper, and the rotation command you are using. For example, if I have a portrait mode document using the US 8.5in by 11in paper, and if \rotateleft{foo} rotates foo 90 degrees counterclockwise, then I can write

     \begin{Landscape}{8.5in}{11in}{\rotateleft}
You can use \LandScape, \Landscape and \Landscape*, rather then the

LandScape, Landscape and Landscape* environments, if you want the rest of the document to be in landscape mode.

If your document is being printed in landscape mode, then these environments switch to portrait mode.

For example, suppose I have a landscape mode document, and I want to switch to portrait mode for the rest of the document, rotating the pages to the “right” with \rotateright. Then I would write

     \Landscape*{11in}{8.5in}{\rotateright}

These environments switch the text width and height, leaving the margins exactly as they were before. It is quite possible that you want to make other changes to the page parameters after switching to landscape mode, but as Lamport points out the LaTEX User’s Guide and Reference Manual, this generally doesn’t work right in the middle of the document. fancybox.sty has a command \UsePageParameters which gets around this. It should be used right after you change the page parameters (and the page parameter changes should come right after the beginning of the landscape environment,

or \clearpage).

Warning: The commands and environments described in this section change LaTEX’s output routine, and may not work with document styles that do the same. Also, bad arguments can cause serious errors with uninformative error messages.

11 Verbatim

If you try to frame some verbatim text by typing

     \fbox{%
       \begin{minipage}{5cm}
         \begin{verbatim}
           \My \Program \Listing
           if { foo } { bar } fi
         \end{verbatim}
       \end{minipage}}

you will get nonsense at best. This is because the argument to \fbox is read before the \begin{verbatim} is processed. But then it is too late for TEX to go back and interpret the contents of the verbatim environment literally rather than as special TEX commands and characters.

One solution is to use the Sbox environment:

     \begin{Sbox}
     \begin{minipage}{5cm}
     \begin{verbatim}
     \My \Program \Listing

     if { foo } { bar } fi
     \end{verbatim}
     \end{minipage}
     \end{Sbox}
     \setlength{\fboxsep}{8pt}
     \fbox{\TheSbox}

                   \My \Program \Listing
                   if { foo } { bar } fi

fancybox.sty also contains a command that “fixes” LaTEX’s LR-box commands for use with verbatim text:

     \VerbBox{cmd}{LR stuff }

This is like

cmd{LR stuff }

but LR stuff can contain verbatim text.^{10} For example:

     \VerbBox{\fbox}{\verb+\foo{bar}+}

                           \foo{bar}

For footnotes, put the command \VerbatimFootnotes in the preamble, and then you can use verbatim commands or environments in the argument of \footnote. This is an optional feature because it might conflict with somebody else’s modification of the footnote system.

If you try to define your own framed verbatim environment with

     \newenvironment{FramedVerb}%
       {\begin{Sbox}\begin{minipage}{5in}\begin{verbatim}}
       {\end{verbatim}\end{minipage}\end{Sbox}
         \setlength{\fboxsep}{8pt}\fbox{\TheSbox}}

10Or other tricks that involve \catcode changes, as occurs with some foreign language macros.

and then type

     \begin{FramedVerb}
       if { foo } { bar } fi
     \end{FramedVerb}

you will again run into trouble because after the \begin{verbatim}, LaTEX is searching for the literal string \end{verbatim} as the end of the verbatim text. It just skips right over the \end{FramedVerb} and may well continue to the end of the file or until it throws up.

fancybox.sty contains some verbatim environments that get around this problem and that have other advantages for LR-boxing verbatim listings, when compared to the standard LaTEX verbatim environment. Admittedly, many of their special features have nothing to do with boxes.

Here are the basic verbatim environments:

Verbatim Works pretty much like LaTEX’s verbatim.

LVerbatim Like Verbatim, but list rather than trivlist is used to display the listing, and so it is indented from the left margin. (This is what I am using for verbatim listings in this document.)

BVerbatim[pos] Produces a box with the same width as the longest verbatim line. The baseline is in the center, unless you include the optional argument [t] for alignment with the top line or [b] for alignment with the bottom line.

VerbatimOut{file} Writes the verbatim text to file.

SaveVerbatim{cmd} Saves the verbatim text as cmd. cmd is defined globally, without checking whether cmd is already defined. Use obviously innocuous names like \MyTemp.

Important: For any of these verbatim environments, or new verbatim environments you define yourself (see below), nothing should come after

\begin{Verbatim} or before \end{Verbatim} on the same line — not even

spaces!^{11} If you put something after \begin{Verbatim} on the same line, it is simply ignored. However, if you put something after \end{Verbatim} on the same line, or if you misspell \end{Verbatim}, you will get an error such as

11If you need to allow something to come before \end{Verbatim}, then you have two options:

• Put the command \AltGetVerbatim in the preamble. This switches to a scheme

     ! File ended while scanning use of \Verbatim.

and the document will end at that point.

You can define new verbatim environments using \newenvironment. You just have to start the definition with

     \VerbatimEnvironment

For example, here is the framed verbatim environment we tried earlier:

     \newenvironment{FramedVerb}%
       {\VerbatimEnvironment
         \begin{Sbox}\begin{minipage}{5cm}\begin{Verbatim}}%
       {\end{Verbatim}\end{minipage}\end{Sbox}
         \setlength{\fboxsep}{8pt}\fbox{\TheSbox}}

Let’s give it a try:

     \begin{FramedVerb}
     if { foo } { bar } fi
     \end{FramedVerb}

                    if { foo } { bar } fi

Here are three commands for inputting a whole file verbatim. The file must end with a new line.

where anything preceding \end{Verbatim} is simply ignored. This can cause prob-

lems if you do really weird things with active characters or other commands within

the verbatim environment (e.g., active conditionals that are not balanced within a

line of verbatim text), but in this case you are probably a good enough hacker to

use the next option.

• \EndVerbatimTokens is a token register that you can set to the tokens that should

precede \end{Verbatim} on the same line, with their verbatim \catcode’s.

\VerbatimInput{file} Like \Verbatim.
\LVerbatimInput{file} Like \LVerbatim.
\BVerbatimInput[pos]{file} Like \BVerbatim.

Here are three commands for making use of verbatim text that has been saved to a command:

\UseVerbatim{cmd} Like \Verbatim.
\LUseVerbatim{cmd} Like \LVerbatim.
\BUseVerbatim[pos]{cmd} Like \BVerbatim.

The SaveVerbatim environment and the \UseVerbatim commands are useful for including verbatim text in the argument of \marginpar, \fbox and other commands. For example, here is another way to define the FramedVerb environment:

     \newenvironment{FramedVerb}%
       {\VerbatimEnvironment
         \begin{SaveVerbatim}{\MyTemp}}%
       {\end{SaveVerbatim}%
         \setlength{\fboxsep}{8pt}%
         \fbox{\begin{minipage}{5cm}\UseVerbatim{\MyTemp}
           \end{minipage}}}

Here are some verbatim commands for short-pieces of (in-line) verbatim text:

\Verbchar literal char Like LaTEX’s \verb command, but it will complain if it encounters a new line in literal.^{12} For example:

         The main use for the \Verb+SaveVerbatim+
         environment and the \Verb+\UseVerbatim+
         commands is to include $\ldots$

The main use for the SaveVerbatim environment and the \UseVerbatim commands is to include … \UseVerb{cmd} Like \UseVerbatim, but without any particular format-

12Be careful that your word processing does not insert one for you.

ting. It is intended for including short pieces of literal text saved with \SaveVerb (below).^{13}

\SaveVerb[whatever]{cmd}char literal char This is like \Verb, but it saves literal as cmd, and then returns to the optional argument whatever. Like the SaveVerbatim environment, it defines cmd globally without checking whether cmd is already defined. Without the optional argument, the most common use is for including verbatim text in a \marginpar, \section or other command argument.

The optional argument can be used for special tricks. For example, all the listings of commands in this documentation use \vitem in a description environment, where \vitem is defined by:^{14}

    \newcommand{\vitem}%
      {\SaveVerb[{\item[\UseVerb{\MyTemp}]}]{\MyTemp}}

Whereas

    \item[\Verb"\foo"]

would not work because after \item reads its argument it is too late to interpret \foo literally,

    \vitem"foo"

does work because it is equivalent to

    \SaveVerb{\MyTemp}"foo"\item[\UseVerb{\MyTemp}]

These environments and commands use various parameters that make it easy to customize their behavior. However, until you want to find the need for such customization, you might as well ignore the rest of this section.

Internally, fancybox.sty separates the reading and formatting of verbatim text. Most of the environments and commands perform both functions, but

SaveVerbatim and \SaveVerb only read the text, while UseVerbatim (and

company) and \UseVerb only format the text. VerbatimOut gets special treatment. The parameters that apply to each class of verbatim environment or command is listed in Table 3.

13But it can also be used for multiple lines saved with the SaveVerbatim environment if you want to do the formatting yourself. E.g., try this in a tabbing environment with \VerbatimTab appropriately defined.

14The braces enclosing the optional argument of \SaveVerb prevent the ] inside the argument from being mistaken for the end of the argument.

Where  What

Environments that format

\VerbatimSpace
\VerbatimTab
\VerbatimFont
\VerbatimFuzz
\EveryVerbatimLine
\EveryVerbatim
\ThisVerb

Environments that read

\EveryVerbatimCodes
\ThisVerbCodes

\Verb and \UseVerb

\VerbSpace \VerbTab \VerbFont \EveryVerb \ThisVerb

\Verb and \SaveVerb

\EveryVerbCodes
\ThisVerbCodes

VerbatimOut

\EveryVerbOutCodes
\ThisVerbCodes
\EveryVerbOutLine
\ThisVerb

Table 3: Parameters for the verbatim environments and commands.

All the parameters, including \VerbatimFuzz, are ordinary commands, and should be changed with \renewcommand.

Here is a description of each of the parameters for environments that format the verbatim text:

\VerbatimSpace The insertion text for spaces. The default is \ , which produces a blank space. Change it to \ttspace to get ␣.

\VerbatimTab The insertion text for tabs. The default is

    \ \ \ \ \ \ \ \

\VerbatimFont The font to use for verbatim text. The default is \tt

\VerbatimFuzz This is the amount by which lines can be too long in a Verbatim or LVerbatim environment before you get overfull \hbox warnings. This threshold is usually .1pt, but the default definition of \VerbatimFuzz is 2pt because verbatim lines won’t break and are therefore often too long.

\EveryVerbatimLine This is inserted at the beginning of each line of verbatim environments or verbatim files. By default it does nothing. I like to indent each line in the verbatim environment in the input file by 2 spaces, so I define

    \renewcommand{\EveryVerbatimLine}[2]{}

to eat those spaces. (But I have to remember to put in two spaces or space markers for blank lines too.) You might also use it to number the lines. For example:

    \newcounter{VerbLineNo}
    \renewcommand{\EveryVerbatimLine}%
      {\makebox[10pt][r]{%
         \stepcounter{VerbLineNo}%
         \tiny\rm\arabic{VerbLineNo}}%
      \hspace{10pt}}
    \renewcommand{\EveryVerbatim}%
      {\setcounter{VerbLineNo}{0}}
    \begin{SaveVerbatim}{\MyTemp}
    \setlength{\fboxsep}{15pt}
    \setlength{\mylength}{\linewidth}
    \end{SaveVerbatim}
    \fbox{\BUseVerbatim{\MyTemp}}

                 1 \setlength{\fboxsep}{15pt}
                 2 \setlength{\mylength}{\linewidth}

\EveryVerbatim Whatever else you want to say before formatting the verbatim text. By default, it does nothing.

\ThisVerb This is executed before any of the commands above, and then its value is cleared. Use this to customize a single verbatim formatting environment.

Here is a description of the parameters for environments that read the verbatim text:

\EveryVerbatimCodes This command is inserted just before reading the verbatim text. Use it to play with \catcode’s (see the TEXbook). For example, I might type

         \renewcommand{\EveryVerbatimCodes}{\catcode`\"=14}

if I want to use " as a comment character in verbatim text.^{15}

\ThisVerbCodes This command is executed before \EveryVerbCodes, and then it is cleared. Use this to fool with the \catcode‘s of a single verbatim environment.

The parameters for \Verb, \UseVerb and \SaveVerb and the VerbatimOut

environment are analogous to the similar commands for other environments.

Here is an example of the use of \ThisVerb to define a variant of \Verb that uses ␣ to mark spaces:

     \newcommand{\SVerb}{%
       \renewcommand{\ThisVerb}%
         {\renewcommand{\VerbatimSpace}{\ttspace}}%
       \Verb}

15Here is another \catcode trick. We make " a short verbatim command, so that we can say "foo" instead of \Verb"foo":

\def\MyQuote{"}    % \MyQuote is now the character ",

\def\temp{\Verb"}  %    in case I need it.

\catcode`\"=13     % Now " is like a command.

\let"\temp         % Now "foo" is like \Verb"foo"

\def\do{\noexpand\do\noexpand}      % Now " can be used in verbatim

\edef\dospecials{\dospecials\do\"}  % environments anyway.

Finally, without further comment, here are the definitions of the example and example* environments that were used for the examples in this document:

% 1. Save example verbatim to \jobname.tmp,
% 2. Input verbatim with \catcode`\"=14 (" is a comment).
% 3. Input again with \catcode\`"=9 (" is ignored).
\renewcommand{\EveryVerbatimLine}[2]{}
\renewcommand{\EveryVerbOutLine}[2]{}
\newcommand{\BeginExample}{%
  \VerbatimEnvironment\begin{VerbatimOut}{\jobname.tmp}}
\newcommand{\EndExample}{%
  \end{VerbatimOut}%.
  \renewcommand{\EveryVerbatimLine}{}%
  \renewcommand{\EveryVerbatimCodes}{\catcode`\"=14}%
  \LVerbatimInput{\jobname.tmp}%
  \catcode`\"=9}
\newenvironment{example}{\BeginExample}{\EndExample
  \begin{center}\input{\jobname.tmp}\end{center}}
\newenvironment{example*}{\BeginExample}%
  {\EndExample \input{\jobname.tmp}}
\newenvironment{example**}{\BeginExample}%
  {\EndExample \globaldefs=1 \input{\jobname.tmp}}

12 Changes

Version 0.9, November 23, 1992 First release.

Version 0.91, November 25, 1992

1. \shadowsize is now a length, to be set with \setlength or

\addtolength.

2. \fancypage split into \fancypage and \fancyput

Version 0.92, December 20, 1992 1. New \boxput command for superimposing boxes.

2. New \VerbatimFootnotes command.

3. New \VerbBox command for verbatim boxes.

4. New Sbox environment. Makes \beginsbox and \endsbox obsolete, but the latter have been retained for compatibility.

Version 0.93, January 20, 1993

1. New \EndVerbatimTokens for tokens that precede \end{Verbatim}.

2. New \AltGetVerbatim, to allow any tokens to precede

  \end{Verbatim}.

3. Fixed bug in \BVerbatimInput that caused second line to be repeated.

4. New LandScape environment, for rotating pages without rotating headers and footers.

5. Slide frames Oval, oval, shadow and double are defined when seminar.sty v0.93 (and maybe later) is loaded.

Version 1.0, February 10, 1993 1. Fixed bugs in \boxput.

2. New Beqnarray environment.